### v1

Abstract
To tackle the crucial problem of crime, evidence at indoor crime scenes must be analyzed before it becomes contaminated or degraded. Here, as an application of artificial intelligence (AI), computer vision, and robotics, we explore how a blimp could be designed as a kind of "floating camera" to drift over and record evidence with minimal disturbance. In particular, rapid prototyping is used to develop a proof-of-concept to gain insight into what such blimps could do, manually piloted or semi-autonomously. As a result, we show the feasibility of attaching various components to an indoor blimp, and confirm our basic premise, that blimps can sense evidence without producing much wind. Some additional suggestions–regarding mapping, sensing, and path-finding–aim to stimulate the flow of ideas for further exploration.
keywords:
[…]
At the crossroads of applied AI, design, computer vision and robotics, this paper explores how a blimp could be used to non-disruptively analyze indoor crime scenes.
Crime is an important problem, ruining numerous lives: the global cost of violence has been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [1]. Furthermore, current approaches are not always successful, with low "clearance rates" the norm rather than the exception (e.g., 40% of homicides worldwide are not being solved, with worsening rates in various countries) [2]. As voiced in the UN’s 16th sustainable development goal (Peace and Justice), more needs to be done to prevent crime. However, crime often takes place behind closed doors and without witnesses, which requires careful documentation to unravel what happened. In such cases, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene . . . Lost or compromised evidence makes a crime scene investigator’s job harder and can seriously damage investigations. Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111universalclass.com/articles/law/setting-crime-scene-perimeters.htm Evidence can also degrade naturally over time (e.g., DNA due to bacteria, heat, light, moisture, and mold) [3] when investigations are delayed due to under-staffing and under-funding–as in the saying "when seconds count, the police are only minutes away". To preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
The term "blimp" refers to a non-rigid, lighter-than-air aircraft that possesses actuators and sensors within a "gondola" (or "nacelle") to enable flight and monitoring.222https://www.airships.net/dirigible Here we use the term in a relaxed sense, permitting some small rigid structure, in the belief that our ideas could be also generally relevant for other kinds of airships, zeppelins, lighter-than-air unmanned aerial vehicles (LTA-UAVs), dirigibles, balloons, and aerostats. While aerodynes like drones could also be useful for indoor crime scene analysis, we believe that blimps feature some merits such as safety, silence, stability (ability to hover over long periods of time), sustainability (low cost, low energy/emission), soft landing, easy transport when deflated, and/or minimal training/licensing requirements for operators [4]. Conversely, the demerits of blimps, such as slow speed, large size, and sensitivity to wind, might be less problematic indoors in smaller domestic areas that are quick to cover, while still wide enough for humans to move around, and potentially less windy than large outdoor sites.
A challenge is that such blimps intended to operate within difficult human environments in the context of crime would require a complex array of hardware and software capabilities. To tackle this challenge, we adopted a rapid prototyping methodology that emphasizes "flexibility, possibility, and design insights as incubated subjectively through the designer", seeking to "cast our net widely" [5]. Thus, our contribution lies in reporting on some opportunities and limitations revealed during our design experience.
The remainder of the paper is organized as follows: In Section 2, we summarize some previous work on blimps for security and indoor use. Details from our exploration in Section 3 are summarized in Section 4, which also points out some next waypoints on our journey to supporting safe and just societies.
2 Related Work
The basic promise of flying platforms for facilitating crime scene analysis has been explored, by manually piloting (or simulating flight of) a drone to spot mock-up bloodstains, guns, knives, and bodies [6, 7, 8]. We have also previously presented an overview of various tasks such a flying robot could perform at a crime scene [9]. However, Bucknell and Bassindale raised the alarm that downwash (wind from a drone’s propellers) could disturb sensitive evidence [10]. This could be problematic, since fibers and fine particles have been used, in some important but difficult cases, to catch various criminals, such as the "Route 40 Killer" or "Green River Killer".333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In the end, Bucknell and Bassindale concluded that drones should be flown at high heights [10]. This was a consequence of how this investigation was conducted, which involved piloting a Parrot AR.Drone 2.0 at various heights. However, some wind could still be generated, and a drone might not be able to distance itself sufficiently when flying under low ceilings or detecting high-up evidence on a table or shelf. Thus, it seemed likely to us that less wind might be generated by a blimp slowly drifting or hovering overhead without using propellers, which is our focus here.
More generally, the use of blimps for security applications has been considered previously: For example, blimps could be used by the police to monitor and deter crime [11].555http://www.reuters.com/article/2011/01/16/us-crime-blimp-utah-idUSTRE70F1DJ20110116 Saiki et al. proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [12]. As well, the Ukrainian Armed Forces are also allegedly using tethered aerostats with infrared sensing from Aerobavovna and Kvertus that can help with surveillance and communication, act as signal relays, and launch drones when an enemy drone is detected [13]. The same article points out that China and the United States have also proposed aerostats that can launch drone swarms–an interesting idea also for crime scene analysis.
Various companies have also focused on designing small indoor blimps for entertainment: For example, a patent from 1994 (US5429542A) proposes a helium-filled remote-controlled saucer toy, and various toys have been sold, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo released many imaginative creations, such as a floating manta ray called in 2007, 888https://www.youtube.com/watch?v=c3-wIICjAhE a jellyfish called the Festo AirJelly in 2008, 999https://www.youtube.com/watch?v=divLsTtA5vk a penguin called the AirPenguin in 2009, 101010https://www.youtube.com/watch?v=jPGgl5VH5go a moving origami-style blimp called the SmartInversion in 2012, 111111https://www.youtube.com/watch?v=BDiTicX8HRU a swarm of dancing spherical blimps with LEDs called the eMotionSpheres in 2014, 121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and a spherical blimp that can pick up a bottle and hand it to a person, called FreeMotionHandling, in 2016. 131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
Papers have also explored the design of novel kinds of indoor blimps for various applications. For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people; in addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, with wings or propellers [14]. 141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In 2015, St-Onge et al. also made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [15]. As well, in 2019, Yao et al. built a blimp to safely follow people indoors, detecting faces and two hand gestures via an RGB camera, and communicating via visual flight patterns and an LED display [16]. Furthermore, in 2019, Ferdous et al. built a blimp with low-current components, including an IR sensor and IMU, to enable usage over many hours [17]. In 2020, Seguin et al. used a convolutional neural network (CNN) to estimate the horizontal position and heading of a blimp from RGB video, along with a LiDAR to estimate height [18]. Recently, in 2024, Pham et al. described a bio-inspired blimp with fins and an onboard inertial measurement unit (IMU) [19]. Additionally, in 2024, Huang et al. designed an interesting 36" blimp with an OpenMV camera that can conduct remote measurement of important vital signs such as heart rate, respiration, or blood pressure via remote photoplethysmography; the blimp detects people with YOLO, approaches via PID control, and transmits data from stably detected foreheads [20]. Furthermore, in 2025, Xu et al. described a fish-shaped flapping-wing blimp named Cuddle-Fish that users enjoyed touching [21]. While offering insights, none of these studies considered the unique challenges of crime scene analysis by a blimp.
With regard to functions for movement, documentation, and sensing, much prior work with blimps seems to have focused on control algorithms to deal with wind for particular embodiments [22]–suggesting the usefulness of checking how easy it is to control the blimp manually. As well, videos taken by a blimp’s camera would require time to watch; given that various previous work has explored the use of 3D mapping of crime scenes [23], it seemed useful if the video could be used to generate a 3D model that can be immediately inspected. Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar, or WiFi signals). For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [24], or drugs.
Looking further into the future, it could also be useful to check the feasibility of automatically detecting and classifying evidence, and how a blimp could plan itself how to move around to acquire data. For example, we have previously used YOLO, to detect suspicious objects [25]. Such tools could be used to gain insight into an important kind of evidence, blood. Bloodstains are fundamentally divided into three main kinds: passive drops falling due to gravity, active blood (e.g., flying from a victim’s body or from a weapon), or transfer stains (e.g., smears from a bloody hand or shoe walking through blood).151515https://www.forensicsciencesimplified.org/blood/principles.html/ While Bergman et al. recently described using a CNN to discriminate between passive drip vs. active spatter bloodstains [26], we are not aware of work that has tackled classifying between all three classes (possibly due to the sensitivity and paucity of publicly available crime data).
As well, in addition to making sense of what it "sees", an autonomous blimp would also be required to itself plan how to move to acquire data. The problem of how a robot should move to cover an area is known as coverage path planning (CPP). An NP-hard problem related to the Traveling Salesman Problem, CPP is relevant to various applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [27]. For example, typical simplified algorithms for a vacuum cleaner robot include snaking (also known as zigzag, snake, or s-path algorithm–typically combined with a random walk), as well as random motion, and spiraling [28]. Snaking was found to be comparatively efficient for achieving high coverage faster, with less unneeded passes–at the cost of more turning and potential weakness to wheel or sensor errors that could occur in cheaper robots. In more complex scenarios where optimized algorithms are desired, the state of the art involves deep reinforcement learning (RL; e.g., for vacuum cleaners [29]). Thus, various functions could be useful, which invited exploration.
3 Methods
To start exploring potential manual and automated capabilities, a basic platform was first required. With the aim of facilitating replication, cost-effective, standard-sized, off-the-shelf components and freely available software tools were used, as described below.
3.1 Set-up
3.1.1 Hardware
Various electronics components where used, such as those summarized in Table 1.
| Component | Weight | Current |
| Raspberry Pi Zero 2 W | 10 g | 460mA161616https://andrejacobs.org/b24/electronics/raspberry-pi-zero-2-w-temperature-and-power-consumption/,171717https://picockpit.com/raspberry-pi/everything-about-raspberry-pi-zero-2-w/ |
| ESP32-Cam | 10 g | 180- 310mA181818https://docs.sunfounder.com/projects/galaxy-rvr/en/latest/hardware/cpn_esp_32_cam.html,191919https://blog.arducam.com/esp32-machine-vision-learning-guide/ |
| TF-Luna | 5 g | 70mA (x 3 = 210mA)202020https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor |
| MLX90640 | 3.5 g | 20mA212121https://www.adafruit.com/product/4407 |
| Lipo Shim | 2.6 g | (15uA quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
Computing. To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5cm × 3.0cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to facilitate user control of the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM was used.
Sensors. RGB Camera. A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera. 242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera, which although costing more, features more onboard processing capabilities.)
Lidars. To detect distances in front, to the side, and below the blimp (e.g., to enable wall-following or mapping), three TF-Lunas were used. TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on a 850nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.252525https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, various useful alternatives existed, like the VL53l1X sensor with 4 m range, but the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
Thermal camera. Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached a MLX90640 board. MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with 110 x 70 field of view composed of a 24 x 32 array of IR thermal sensors that detect from -40 C to 300 C with an accuracy of 2C at a frame rate of 2 Hz.262626https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array
Power. To power the microcomputer and sensors, a "LiPo Shim" was used, which can provide a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).272727https://shop.pimoroni.com/products/lipo-shim With our two small batteries (1000 mAh or 700 mAh), operation was approximately 20 minutes. As well, a separate 3 V CR2 battery was used to power the actuators.
Actuation and Body. To achieve motion without "reinventing the wheel", we modified an existing toy, comprising three DC motors attached to propellers and a radio receiver/transmitter. The propellers allow left and right rotation, as well as upward, downward, forward, and backward translation. To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.282828https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total). (Given the weight of helium and air (0.18 kg/m3 vs. 1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)292929https://specialtygases.messergroup.com/documents/20136/3998363/Balloon%2BHelium.pdf
During our design exploration, one observation that stood out was how tricky it can be to design small blimps with many sensors: Size. First, we found with a miniature toy unit and regular sized latex balloon that virtually no payload could be accommodated and small balloons lost helium quickly. Next, we tried using two balloons with a larger toy unit, but even though each individual sensor could be fairly light, the weight of busses, cables, tape and boards to connect and affix sensors quickly added up. Furthermore, much current was required for the microcomputer, sensors, and actuators, which in turn called for including an appropriately large battery. As well, to avoid voltage spikes from the motors resetting the microcomputer, separate power supplies were used. Thus, in the spirit of easing such restrictions and enabling freer exploration, our latest version became quite large, with three balloons (total height approximately 150 m with gondola and ballast pouch). Awkwardness. Some initial attempts to use single and double sided tape and velcro to connect the gondola and balloons didn’t provide good results, and aligning bulky components seeking to rise or fall was difficult. The gondola and balloons sometimes split, with heavy tape and velcro adhering just to one surface and ripping away material, creating unclean surfaces. Previously, working with the "Angel" blimp, we had attached hooks to a table and each balloon, but this time we wished avoid adding such weight. Instead, individual balloons were temporarily stowed under tables and desks to immobilize them. When combining balloons, we tried at one point using a low ceiling (2.3 m) in a corridor. However, this was a public space with people frequently needing to pass, and once, after returning from a quick errand, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart. Thereafter, we propped the blimp against the bottom of a projector in our lab, which had a high ceiling (3 m+). We tied strings through each balloon’s tail section (where helium is inserted) and holes in the projector shelf, as well as to chairs below. To connect the balloons more securely, we modified some wall hooks to attach capped carbon fiber rods running up along both sides of the blimp. Even then, the top balloon sometimes flew off, so we attached set screws to both the bottom and top of the blimp to secure the rods firmly. (While we could have used a bigger balloon instead of using multiple smaller balloons, we believe this modularity is useful for exploratory design, since we can shrink or expand freely, also without wasting helium.) As well, to prevent the gondola from falling off, we made a plastic base that was firmly taped to the bottom balloon and further secured with threads going to the side hooks; then bolts and nuts were used to attach and remove the gondola. Another challenge was that the sensors had to be below the propellers, so rods were used to suspend a "second" small gondola from the first, built from light cardboard. Additionally, ballast was used to achieve neutral buoyancy in air, but we noticed that recalibration was required whenever we switched out different-sized batteries (1000 mAh or 700 mAh). As well, when testing the propellers of the gondola on a table, we were surprised when even powering motors for a brief instant was even to cause the entire gondola to spin into our laptop, causing a propeller to break; thereafter we took care to firmly tape down the gondola whenever testing it.
3.1.2 Software
Software was also required. Server and client programs were written to send LiDAR, RGB, and thermal data, as well as control signals, between the gondola and an external laptop. TigerVNC was used to remotely develop the software on the Raspberry Pi.303030https://tigervnc.org/ As well, we built a user interface with color-coded functions in Python and OpenCV,313131https://opencv.org/ shown in Fig. 3. While we created Arduino code to control the blimp in various ways, to avoid potential problems if communication broke down, we ended up using commands that initiate short bursts rather than continuous motion.
One observation when developing software was how little information we could find on how to set up a reasonably realistic mock-up crime scene for testing; e.g., what kind of evidence is typical and where it is usually found. In addition to the challenge that crime classification schemes differ regionally, in line with the diversity of criminal law systems, Google search hits for random crime scene generation seemed to bring up text summaries of crimes intended to help stimulate writers (e.g., "Murder Scene Generator"323232https://perchance.org/znevive9bn). We could also not find any datasets online of crime scene sketches. One reason for the lack of crime data could be due to a view that such data are sensitive and private, for victims and their families, and could also spark copy-cat crimes. While one or two mock-up sketches were available (e.g., Galanakis et al. made available a scan of one mock-up crime scene [30], and our partners at the Swedish Police also shared a scan of a mock-up crime scene at NFC Villa), as the saying goes, no two crime scenes are the same. Thus, to get started, we built two simplified programs to simulate how evidence might be distributed at a crime scene, as shown in Fig. 4. The first general program allows a user to select a floor plan and a kind of crime (homicide, assault, burglary, or arson), then randomly computes which kinds of evidence are found and where. For initial floor plans, we used (1) a LiDAR scan of the National Forensics Center (NFC) villa, used by our partners at the Swedish Police to train, and (2) a virtual depiction of Halmstad Intelligent Home (HINT), a 50 m2 mock-up sensorized living environment at our university. All probability parameters can be assigned or tweaked by the user. The second program, intended for initial testing, simply creates a single "heap" of evidence, taking a list of evidence and using it to fill a grid. Large evidence comprising more than one cell is filled in first, then smaller evidence.
3.2 Exploring Possibilities for Indoor CSA Blimps
With the developed platform, based on the literature, we speculated that it could be interesting to check how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used. Furthermore, looking toward the future, we wished to check automatic detection, classification of bloodstains, and how a blimp could plan itself how to move around to acquire data. For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, and one small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye–E120) to simulate passive, active, and transfer stains.
3.2.1 Disturbance
To test the degree to which sensitive evidence might be disturbed by a blimp, we placed a Mastech MS6252A digital anemometer on the ground beside the mock-up evidence. As a result, the wind sensor reading was 0.0 m/s even when passing the blimp even quite close to the evidence ( 20 cm). For comparison, passing a typical small drone, the DJI Ryze Tello, over the evidence at 1.2 m resulted in a wind speed of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
3.2.2 Sensing while drifting
Next, we checked on the blimp’s ability to move and sense evidence. The blimp was controlled manually to obtain 10 data samples. In each trial, a random ordering of the evidence was calculated using our Python program, then the physical evidence was manually placed by a human. As a result, on average, the blimp captured 5.2 of 7 objects (74.3%) in its view. The reason for missing some evidence sometimes was because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
3.2.3 3D Mapping
For easy viewing, we wished to explore converting video from the blimp to a 3D model. To test the waters, a video from the blimp was passed to the free version of 3DF Zephyr to create a 3D textured mesh. The result is shown in the Fig. 5(a). Some rough areas existed, e.g., on the mannequin "corpse" to the left-hand side of the image), since the free version of this software only uses the first 50 photos extracted from the video. (Due to this limitation, we also sped up the video to ensure the whole scene would be included.) However, we believe that the time required to generate a 3D map (approximately 3 minutes) was promising.
3.2.4 Thermal Sensing
In each trial controlling the blimp, we experimented with also touching a knife and the shoes directly beforehand. As a result, we noted that some warm areas could be seen in the thermal data, as in Fig. 5(b), but in general, the warmed objects (especially the smaller knife) seemed hard to distinguish due to our sensor’s low resolution. As well, a checkerboard pattern appears due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later). Thus, we believe that a higher resolution sensor with different update process would be required some day in the future for identifying recently-touched, non-metallic objects. (The MLX90640 could however still serve some other use, such as detecting investigators to follow, or hidden persons).
3.2.5 Object detection
It could save time if the blimp could be automated to perform important tasks like detection. This could then be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention to more thoroughly investigate certain areas. As an initial check, we tried inputting an image taken by the blimp to a pretrained generic YOLO detection model (YOLOv8l)333333https://docs.ultralytics.com/ and ChatGPT.343434hhttps://chatgpt.com The first result in Fig. 5 (c) shows that while one knife is detected correctly, a bloody handprint is mistaken as a horse, and other objects are not detected; these errors could be due to low resolution or slightly dark lighting (approximately 300 lux in our lab when the image was taken). ChatGPT, however, correctly detected all objects. This seemed promising to us, and suggested using either an adapted version of YOLO with sufficient data for guns and blood, or a foundation model.
3.2.6 Bloodstain Classification
To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence heap. First we color-picked red, then assessed the area and eccentricity of an ellipse fitted to detected contours of reasonable size. We estimated that transfer patterns from a hand might have high area, passive drops should have low eccentricity, and active drops high eccentricity, as in Fig. 5(d). This resulted in an classification accuracy of 74.1% (20 of 27 detected blood contours being correctly classified). Adding in false positives from color-picking (e.g., detected in the red tags on the shoes), the overall accuracy was 69.0% (20 of 29), which seemed promising for initial prototyping.
3.2.7 Coverage path planning for crime scenes
An automated blimp could move without piloting. Given that a full solution to this complex problem wouldn’t be possible within the scope of the current paper, we brainstormed to identify some variables of interest and formulate the general problem. We believe the blimp’s path planning strategy should depend on:
How precise blimp’s movements are–if imprecise, the blimp should stay farther away to avoid crashing, and more overlap might be required.
How accurate the blimp’s sensors are–if inaccurate, the blimp should stay closer to the evidence to achieve proper documentation.
How urgent the task is–if the blimp is needed elsewhere, a rapid path could be calculated; else in an important case, more time can be taken.
If the blimp should also gather evidence–then it should probably be close.
How much detail is desired–if the blimp should see under obstacles like tables and beds.
In general, however, we believe that indoor blimp crime scene analysis is characterized by the following requirements:
Some speed is desired, given that evidence can become contaminated or degraded over time.
Some overlap is desired for 3D mapping; much (unneeded) overlap however would be undesirable since excess motion could increase the risk of disrupted evidence.
While human domestic environments can be highly complex (e.g., ceiling obstacles like lamps, fans, alarms, or chandeliers, and ground obstacles like chairs, clothes, etc. and L-shaped rooms or hallways and walk-in closets), rooms tend to often be rectangular.
The blimp should try to restrict turning and vertical changes to avoid using up its battery. For example, the blimp can fly over low objects (e.g., to see on tables, beds)–although, compared to an outdoors scene, height is more constrained.
People in the room might intentionally or unintentionally obstruct the robot.
The blimp could be acting as part of a swarm.
The environment might be known (e.g., from floor plans) or unknown.
Based on this, both simplified and more complex solutions seemed possible. For simplicity, we propose that a snaking algorithm combined with a random walk and/or wall-following could be a starting point for simple testing, that is intuitive and easy to implement and carry out. For maximum detail, the blimp could calculate a path where it flies just above surfaces; else for efficiency, the blimp could calculate a space below any ceiling obstacles and above any ground obstacles. As well, the blimp could scan from high-up first, then return to any areas that seem interesting (e.g., where evidence has been discovered). Alternatively, a deep RL algorithm could be used to document in a human-like way (e.g., based on data of human investigators): to achieve high efficiency (reduce flight time, avoiding redundant re-coverage), improve map quality (ensure good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), and be room-agnostic (function successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnostic (different blimps could use the same process to learn how to move well). A cost might be more difficulty in the implementation (complex equations, much training data required, tuning), and loss of explainability.
4 Discussion
In summary, we explored a new design idea, of using a blimp to record sensitive evidence at an indoor crime scene. To our best knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, or three LiDARs, or a minicomputer (rather than a microcontroller), etc., and various hardware challenges were discussed. Some software was also developed to simulate crime scenes. As a result of our exploration, we found some support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum. As well, manually piloting was possible but imperfect (sensing 74.3% of evidence). A 3D map was generated within a few minutes using free software. Some recommendations for thermal sensing were given. Toward automating crime scene analysis, all objects in the blimp’s camera view were detected correctly by ChatGPT. As well, the three main kinds of bloodstain were detected and classified with an accuracy of 74.1% via color-picking and consideration of contour area and eccentricity. Finally, we discussed path-planning considerations and strategies. As well, a video and code have been made available.353535https://youtu.be/5ep9UeChn68,363636https://github.com/martincooney/Crime_Blimp
4.1 Limitations and Future Work
The current study is limited by its exploratory design approach, working with mock-up evidence in a lab. The blimp was large, with component weights unoptimized; surface mount/low power components (like a Raspberry Pi Pico instead of Zero 2 W), or Dist-YOLO in place of LiDARs [31], to reduce weight and size. As well, we used a remote laptop, which, although common in the related studies noted, would not be secure at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications. Our crime scene generating software is currently based on guesses and should be trained with real data to enable more realistic outputs. To help blimp operators, control algorithms should be developed. Other applications of heat sensing, and other miniature thermal cameras, should be tested. A more comprehensive study should be done on identifying and detecting the various kinds of evidence that could be important. For classifying bloodstains, only a single image was used. A reasonable number of (real) training and test data should be used in conjunction with a more advanced, machine-learning approach to achieve better results. As well, real composites of various kinds of bloodstains should be analyzed to infer the story of what happened, and further analysis can reveal, e.g., if a criminal is right or left-handed, the type of weapon used, what made a transfer stain (e.g., a hand or a shoe), which wounds were first, if death was immediate or not, and the number of blows, etc. Furthermore, path planning should be implemented and tested on our prototype.
As well, our focus was limited to sensitive crime evidence, whereas remote-monitoring blimps could be useful for other inspections of sensitive materials or even for helping to reduce the number of investigators ambushed and killed at crime scenes. By investigating these and other related directions, our aim is that opportunities could be found to contribute to a peaceful and happy society for all.
Acknowledgements.
We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as advice from the Swedish Police Authority. The initial idea for this paper stemmed from a conversation with Mikael Lilja (Superintendent, Police Region Öst).References
- Hoeffler [2017] A. Hoeffler, What are the costs of violence?, Politics, Philosophy & Economics 16 (2017) 422–445.
- Sturup et al. [2015] J. Sturup, D. Karlberg, M. Kristiansson, Unsolved homicides in sweden: A population-based study of 264 homicides, Forensic science international 257 (2015) 106–113.
- Benson et al. [2003] B. L. Benson, W. David, P. A. May, K. Harries, A. M. Rosenthal, A. Lorie, M. L. D. Fridell, G. Fisher-Stewart, J. Pedro, T. M. Saavedra, et al., Without a trace? advances in detecting trace evidence, National Institute of Justice 249 (2003).
- Rozhok [2023] A. Rozhok, Modeling of Innovative Lighter-than-Air UAV for Logistics, Surveillance, and Rescue Operations, Ph.D. thesis, Università degli studi di Genova, 2023.
- Zamfirescu-Pereira et al. [2021] J. Zamfirescu-Pereira, D. Sirkin, D. Goedicke, R. Lc, N. Friedman, I. Mandel, N. Martelaro, W. Ju, Fake it to make it: Exploratory prototyping in hri, in: Companion of the 2021 ACM/IEEE international conference on human-robot interaction, 2021, pp. 19–28.
- Urbanová et al. [2017] P. Urbanová, M. Jurda, T. Vojtíšek, J. Krajsa, Using drone-mounted cameras for on-site body documentation: 3d mapping and active survey, Forensic science international 281 (2017) 52–62.
- Georgiou et al. [2022] A. Georgiou, P. Masters, S. Johnson, L. Feetham, Uav-assisted real-time evidence detection in outdoor crime scene investigations, Journal of forensic sciences 67 (2022) 1221–1232.
- Araujo et al. [2019] P. Araujo, J. Fontinele, L. Oliveira, Multi-perspective object detection for remote criminal analysis using drones, IEEE Geoscience and Remote Sensing Letters 17 (2019) 1283–1286.
- Cooney et al. [2025] M. Cooney, S. Ponrajan, F. Alonso-Fernandez, Nano drone-based indoor crime scene analysis, arXiv preprint arXiv:2502.21019 (2025).
- Bucknell and Bassindale [2017] A. Bucknell, T. Bassindale, An investigation into the effect of surveillance drones on textile evidence at crime scenes, Science & justice 57 (2017) 373–375.
- Murphy et al. [2013] L. W. Murphy, C. Calabrese, J. Stanley, C. Crump, The future of drones in america: Law enforcement and privacy considerations, ACLU Statement for the Record for a Senate Judiciary Committee Hearing, American Civil Liberties Union 20 (2013).
- Saiki et al. [2011] H. Saiki, T. Fukao, T. Urakubo, T. Kohno, A path following control method under wind disturbances for outdoor blimp robots, in: 2011 IEEE/SICE International Symposium on System Integration (SII), IEEE, 2011, pp. 978–984.
- Trevithick [2025] J. Trevithick, Balloon-launched drone to intercept long range kamikaze drones emerges in ukraine, The War Zone (TMZ) (2025). URL: https://bmva-archive.org.uk/bmvc/2017/papers/paper188/index.html.
- Cooney et al. [2012] M. Cooney, F. Zanlungo, S. Nishio, H. Ishiguro, Designing a flying humanoid robot (fhr): effects of flight on interactive communication, in: 2012 IEEE RO-MAN: The 21st IEEE International Symposium on Robot and Human Interactive Communication, IEEE, 2012, pp. 364–371.
- St-Onge et al. [2015] D. St-Onge, C. Gosselin, N. Reeves, Dynamic modelling and control of a cubic flying blimp using external motion capture, Proceedings of the Institution of Mechanical Engineers, Part I: Journal of Systems and Control Engineering 229 (2015) 970–982.
- Yao et al. [2019] N.-s. Yao, Q.-y. Tao, W.-y. Liu, Z. Liu, Y. Tian, P.-y. Wang, T. Li, F. Zhang, Autonomous flying blimp interaction with human in an indoor space, Frontiers of Information Technology & Electronic Engineering 20 (2019) 45–59.
- Ferdous et al. [2019] S. U. Ferdous, A. Mohammadi, S. Lakshmanan, Developing a low-cost autonomous blimp with a reduced number of actuators, in: Unmanned Systems Technology XXI, volume 11021, SPIE, 2019, pp. 73–80.
- Seguin et al. [2020] L. Seguin, J. Zheng, A. Li, Q. Tao, F. Zhang, A deep learning approach to localization for navigation on a miniature autonomous blimp, in: 2020 IEEE 16th International Conference on Control & Automation (ICCA), IEEE, 2020, pp. 1130–1136.
- Pham et al. [2024] H. Q. Pham, S. Singh, M. Garratt, S. Ravi, Controlling a bio-inspired miniature blimp using a depth sensing neural-network camera, Bioinspiration & Biomimetics 19 (2024) 024001.
- Huang et al. [2024] H.-W. Huang, J. Chen, P. Rupp, C. Ehmke, P. R. Chai, R. Dhar, I. Ballinger, G. Traverso, Cost-effective blimp for autonomous and continuous vital signs monitoring, in: 2024 IEEE International Conference on Advanced Intelligent Mechatronics (AIM), IEEE, 2024, pp. 1553–1559.
- Xu et al. [2025] M. Xu, J. Shao, Y. Ju, X. Shen, Q. Gao, W. Chen, Q. Zhang, Y. S. Pai, G. Barbareschi, M. Hoppe, et al., Cuddle-fish: Exploring a soft floating robot with flapping wings for physical interactions, arXiv preprint arXiv:2504.01293 (2025).
- Bhat et al. [2024] S. S. Bhat, S. G. Anavatti, M. Garratt, S. Ravi, Review of autonomous outdoor blimps and their applications, Drone Systems and Applications 12 (2024) 1–21.
- Esposito et al. [2023] M. Esposito, F. Sessa, G. Cocimano, P. Zuccarello, S. Roccuzzo, M. Salerno, Advances in technologies in crime scene investigation. diagnostics 2023, 13, 3169. h ps, doi. org/10.3390/diagnostics13203169 Academic Editor: Hiroshi Ikegaya Received 15 (2023).
- Muñoz et al. [2025] J. D. Muñoz, J. Ruiz-Santaquiteria, O. Deniz, G. Bueno, Concealed weapon detection using thermal cameras, Journal of Imaging 11 (2025) 72.
- Cooney et al. [2024] M. Cooney, L. M. W. Klasén, F. Alonso-Fernandez, Designing robots to help women, in: 14th Scandinavian Conference on Artificial Intelligence (SCAI 2024): AI for a better society, June 10-11, 2024, Jönköping, Sweden, IEEE, 2024, pp. 168–177.
- Bergman et al. [2022] T. Bergman, M. Klöden, J. Dreßler, D. Labudde, Automatic classification of bloodstains with deep learning methods, KI-Künstliche Intelligenz 36 (2022) 135–141.
- Jonnarth et al. [2023] A. Jonnarth, J. Zhao, M. Felsberg, Learning coverage paths in unknown environments with deep reinforcement learning, arXiv preprint arXiv:2306.16978 (2023).
- Sörme and Edwards [2018] J. Sörme, T. Edwards, A comparison of path planning algorithms for robotic vacuum cleaners, 2018.
- Moon et al. [2022] W. Moon, B. Park, S. H. Nengroo, T. Kim, D. Har, Path planning of cleaning robot with reinforcement learning, in: 2022 IEEE International Symposium on Robotic and Sensors Environments (ROSE), IEEE, 2022, pp. 1–7.
- Galanakis et al. [2021] G. Galanakis, X. Zabulis, T. Evdaimon, S.-E. Fikenscher, S. Allertseder, T. Tsikrika, S. Vrochidis, A study of 3d digitisation modalities for crime scene investigation, Forensic sciences 1 (2021) 56–85.
- Vajgl et al. [2022] M. Vajgl, P. Hurtik, T. Nejezchleba, Dist-yolo: Fast object detection with distance estimation, Applied Sciences 12 (2022). URL: https://www.mdpi.com/2076-3417/12/3/1354. doi:10.3390/app12031354.